# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v3

@@ Abstract
Human motion generation stands as a significant pursuit in generative computer vision, while achieving long-sequence and efficient motion generation remains challenging. Recent advancements in state space models (SSMs), notably Mamba, have showcased considerable promise in long sequence modeling with an efficient hardware-aware design, which appears to be a promising direction to build motion generation model upon it. Nevertheless, adapting SSMs to motion generation faces hurdles since the lack of a specialized design architecture to model motion sequence. To address these challenges, we propose Motion Mamba, a simple and efficient approach that presents the pioneering motion generation model utilized SSMs. Specifically, we design a Hierarchical Temporal Mamba (HTM) block to process temporal data by ensemble varying numbers of isolated SSM modules across a symmetric U-Net architecture aimed at preserving motion consistency between frames. We also design a Bidirectional Spatial Mamba (BSM) block to bidirectionally process latent poses, to enhance accurate motion generation within a temporal frame. Our proposed method achieves up to 50% FID improvement and up to 4 times faster on the HumanML3D and KIT-ML datasets compared to the previous best diffusion-based method, which demonstrates strong capabilities of high-quality long sequence motion modeling and real-time human motion generation.
@@ 1 Introduction
Human motion generation stands as a holy grail in generative computer vision, holding broad applications in computer animation, game development, and robot manipulation. To emulate human motion effectively, virtual characters must respond to the conditional context, exhibit natural movement, and perform motion accurately. Recent motion generation models are categorized into four main approaches: autoencoder-based [1, 48, 36, 17, 56, 10], utilizing transformers for latent space compression and motion synthesis; GAN-based [30, 20, 4], using discriminators to enhance the realism of generated motions; autoregressive models [24], treating motion sequences as languages with specialized codebooks; and diffusion-based [54, 49, 6], employing denoising steps for motion generation. Challenges vary across methods, with autoencoder models struggling to generate accurate motions from detailed descriptions due to textual information compression, GAN-based models facing training difficulties, especially in conditional tasks, and diffusion-based models relying on complex transformer-based architectures results in inefficient of motion prediction.
@@ 2 Related Works
Human Motion Generation. Generating human motion is a fundamental task of computer vision, essential for various applications like 3D modeling and robot manipulation. Recently, the predominant method of achieving human motion generation, known as the Text-to-Motion task, involves learning a common latent space for both language and motion.
DVGAN [30] create the GAN [11] discriminator by densely validating at each time-scale and perturbing the discriminator input for translation invariance, enabling motion generation and completion. ERD-QV [20] enhances latent representations through two additive modifiers: a time-to-arrival embedding applied universally and an additive scheduled target noise vector used during extended transitions. It further improves transition quality by incorporating a GAN framework with two discriminators operating at different timescales. HP-GAN [4], trained with a modified version of the improved WGAN-GP [16], utilizes a custom loss function designed for human motion prediction. It learns a probability density function of future human poses conditioned on previous poses.
Autoencoders [43, 27] are notable generative models known for their ability to represent data robustly by compressing high-dimensional data into a latent space, which is widely adopted in human motion generation. JL2P [1] uses RNN-based autoencoders [23] to learn a combined representation of language and pose. It restricts the direct mapping from text to motion to a one-to-one relationship. MotionCLIP [48] uses Transformer-based Autoencoders [51] to reconstruct motion while ensuring alignment with the corresponding text label in the CLIP [40] space. This alignment effectively integrates the semantic knowledge from CLIP into the human motion manifold. TEMOS [36] and T2M [17] combine a Transformer-based VAE [26] with a text encoder to generate distribution parameters that work within the VAE latent space. AttT2M [56] and TM2D [10] incorporate a body-part spatio-temporal encoder into VQ-VAE [50] for enhanced learning of a discrete latent space with increased expressiveness.
@@
State Space Models. Recently, state space sequence models (SSMs) [14, 15], drawing inspiration from classical state-space models [25], have emerged as a promising architecture for sequence modeling. Mamba [12] introduces a selective SSM architecture, integrating time-varying parameters into the SSM framework, and proposes a hardware-aware algorithm to facilitate highly efficient training and inference processes. Some research works leverage SSM in computer vision to process 2D data. The 2D SSM [3] introduces an SSM block at the beginning of each transformer block [51, 8]. This approach aims to achieve efficient parameterization, accelerated computation, and a suitable normalization scheme. SGConvNeXt [29] presents a structured global convolution method inspired by ConvNeXt [32], incorporating multi-scale sub-kernels to achieve both parameterization efficiency and effective long sequence modeling. ConvSSM [46] integrates the tensor modeling principles of ConvLSTM [45] with SSMs, elucidating the utilization of parallel scans in convolutional recurrences. This approach enables subquadratic parallelization and rapid autoregressive generation. Vim [57] introduces a bidirectional SSM block [52] for efficient and versatile visual representation learning, achieving performance comparable to established ViT [8] methods. VMamba [31] introduces a Cross-Scan Module (CSM) designed to traverse the spatial domain and transform any non-causal visual image into ordered patch sequences. This approach achieves linear complexity while preserving global receptive fields. There have also been attempts to utilize SSMs to handle higher-dimensional data. Mamba-ND [28] explores various combinations of SSM and different scan directions within the SSM block to adapt Mamba [12] to higher-dimensional tasks. Recent efforts have sought to replace the traditional transformer-based U-Net within the diffusion denoiser with the SSM block, with the aim of enhancing image generation efficiency. DiffuSSM [53] adeptly manages higher resolutions without relying on global compression, thus maintaining detailed image representation throughout the diffusion process.
@@ 3.1 Preliminaries
Selective Structured State Space Sequence Model. SSMs particularly through the contributions of structured state space sequence models (S4) and Mamba, have demonstrated exceptional proficiency in handling long sequences. These models operationalize the mapping of a 1-D function or sequence, $x(t)\in\mathbb{R}\mapsto y(t)\in\mathbb{R}$, through a hidden state $h(t)\in\mathbb{R}^{N}$, employing $\mathbf{A}\in\mathbb{R}^{N\times N}$ as the evolution parameters, $\mathbf{B}\in\mathbb{R}^{N\times 1}$ and $\mathbf{C}\in\mathbb{R}^{1\times N}$ as the projection parameters, respectively. The continuous system dynamics are described by the ordinary differential equation (ODE):
$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t),$ (1)
$y(t)$ $=\mathbf{C}h(t)+\mathbf{D}x(t),$
with $x(t)$ representing a continuous input signal and $y(t)$ a continuous output signal in the time domain.
To adapt these continuous dynamics for practical computation, the S4 and Mamba models employ a discretization process, notably using the zero-order hold (ZOH) method, resulting in a transformation of continuous parameters into discrete ones:
$\mathbf{\overline{A}}$ $=\exp{(\mathbf{\Delta}\mathbf{A})},$ (2)
$\mathbf{\overline{B}}$ $=(\mathbf{\Delta}\mathbf{A})^{-1}(\exp{(\mathbf{\Delta}\mathbf{A})}-\mathbf{I})\cdot\mathbf{\Delta}\mathbf{B}.$
@@ 3.2 Motion Mamba
$K=\{S^{2N-1},S^{2(N-1)-1},\ldots,S^{1}\}.$ (6)
This sequence specifies the number of scans allocated to each layer, in descending order of complexity. For instance, the uppermost encoder layer, $E_{1}$, and the lowermost decoder layer, $D_{N}$, are allocated $S^{2N-1}$ scans, indicating the highest scanning complexity. Conversely, the lowest encoder layer, $E_{N}$, and the uppermost decoder layer, $D_{1}$, are assigned $S^{1}$ scans, reflecting the lowest level of scanning complexity.
$E_{i}(S)=\begin{cases}S^{2N-1}&\text{for }i=1\\ S^{2(N-i)-1}&\text{for }i=2,\ldots,N-1\\ S^{1}&\text{for }i=N\end{cases}$ (7)
$D_{j}(S)=\begin{cases}S^{2N-1}&\text{for }j=N\\ S^{2(N-j)-1}&\text{for }j=N-1,\ldots,2\\ S^{1}&\text{for }j=1\end{cases}$ (8)
@@
Hierarchical Temporal Mamba (HTM) block processes compressed latent representations, denoted as $z$, with the dimensions $(T,B,C)$, of which procedure shown in Algorithm 1. Here, $T$ signifies the temporal dimension, as specified in the Variational AutoEncoder (VAE) framework. Initially, the input $z$ is subjected to a linear projection layer, producing transformed representations $x$ and $z$ with dimension $E$. Our analysis revealed an increased density of motion within the lower-level feature spaces. Consequently, we developed a hierarchical scanning methodology that is executed at various depths of the network. This approach not only accommodates the diverse motion densities, but also significantly reduces computational overhead. This step utilizes a hierarchically structured set of scans, $K=\{S^{2N_{n}-1},S^{2N_{n-1}-1},\ldots,S^{1}\}$, in conjunction with a corresponding series of memory matrices $\{A_{1},\ldots,A_{k}\}$. Each sub-SSM scan first applies a 1-D convolution to $x$, resulting in $x^{\prime}_{o}$. $x^{\prime}_{o}$ is then linearly projected to derive $B_{o}$, $C_{o}$, and $\Delta_{o}$. These projections $B_{o}$, $C_{o}$ use $\Delta_{o}$ to effect transformations in $\overline{A}_{o}$ and $\overline{B}_{o}$, respectively. After executing a sequence of SSM scans $\{SSM_{A_{1},x},SSM_{A_{2},x},\ldots,SSM_{A_{k},x}\}$, a set of outputs $\{O_{1},\ldots,O_{k}\}$ is compiled. This collection is subsequently aggregated via a linear projection to obtain the final output of the HTM block.
@@ 4.3 Comparative Studies
We evaluate our method against the state-of-the-art methods on the HumanML3D [17] and KIT-ML [38] datasets. We train our Motion Mamba with HTM arrangement strategy MM ($\{S^{2N_{n}-1},\ldots,S^{1}\}$), BSM bidirectional block strategy on the latent dimension = 2 with 11 layers. We evaluate our model and previous works with suggested metrics in HumanML3D [17] and calculate 95% confidence interval by repeat evaluation 20 times. The results for the HumanML3D dataset are presented in Table 1. Our model outperforms other methods significantly across various evaluation metrics, including FID, R precision, multi-modal distance, and diversity. For instance, our Motion Mamba outperforms previous best diffusion based motion generation MLD by 40.5% in terms of FID, and up to 10% improvement on R Precision, we aslo obatined best MModality by 3.060. The results for the KIT-ML dataset are presented in Table 2. We have also outperformed other well-established methods in FID and multi-modal distance.
@@ 4.4 Ablation Studies
Long Sequence Motion Generation. The HumanML3D [17] dataset exhibits a long-tailed and right-skewed distribution with a significant proportion of long-sequence human motions, as shown in Figure 3. We suggest previous study overlooked the challenges in the long-sequence generation problem. Thus, we introduce a new dataset variant, HumanML3D-LS, comprising motion sequences longer than 190 frames extracted from the original test set. This addition allows us to showcase our capability in generating long-sequence motions. Subsequently, we evaluate the performance of our method on HumanML3D-LS and compare it with other diffusion-based motion generation approaches. The comparative results are presented in Table 3. Motion Mamba by leverage the benefits on long-range dependency modeling make it well suitable for long sequence motion generation.
Hierarchical Design with HTM. In our ablation studies, we observed a slight improvement upon reversing the scan order from a lower to a higher level, specifically transitioning from MM $\{S^{1},\ldots,S^{N}\}$ to MM $\{S^{N},\ldots,S^{1}\}$. This enhancement suggests a correlation with the increase in temporal motion density within the lower-level feature spaces. Furthermore, to achieve the optimal result, we introduce the hierarchical design to arrange the scanning frequency, resulting in the sequence MM $\{S^{2N_{n}-1},\ldots,S^{1}\}$. This expansion in the number of scans led to a performance increase. We attribute this enhancement to the observation that individual selective scan operations significantly reduce the parameter count, especially when compared to the parameter-intensive constructs of self-attention and feedforward network blocks prevalent in transformer architectures.
@@
MM ($\{S^{1},\ldots,S^{N}\}$) $0.673^{\pm.003}$ $1.278^{\pm.012}$ $3.802^{\pm.041}$ $8.678^{\pm.096}$ $3.127^{\pm.024}$
MM ($\{S^{N},\ldots,S^{1}\}$) $0.738^{\pm.002}$ $0.962^{\pm.011}$ $3.433^{\pm.003}$ $9.180^{\pm.071}$ $2.723^{\pm.033}$
MM ($\{S^{1},\ldots,S^{2N_{n-1}}\}$) $0.698^{\pm.002}$ $0.856^{\pm.008}$ $3.624^{\pm.037}$ $9.229^{\pm.067}$ $2.826^{\pm.017}$
MM ($\{S^{2N_{n}-1},\ldots,S^{1}\}$) $0.792^{\pm.002}$ $0.281^{\pm.009}$ $3.060^{\pm.058}$ $9.871^{\pm.084}$ $2.294^{\pm.058}$

### v4

@@ Abstract
Human motion generation stands as a significant pursuit in generative computer vision, while achieving long-sequence and efficient motion generation remains challenging. Recent advancements in state space models (SSMs), notably Mamba, have showcased considerable promise in long sequence modeling with an efficient hardware-aware design, which appears to be a significant direction upon building motion generation model. Nevertheless, adapting SSMs to motion generation faces hurdles since the lack of a specialized design architecture to model motion sequence. To address these challenges, we propose Motion Mamba, a simple yet efficient approach that presents the pioneering motion generation model utilized SSMs. Specifically, we design a Hierarchical Temporal Mamba (HTM) block to process temporal data by ensemble varying numbers of isolated SSM modules across a symmetric U-Net architecture aimed at preserving motion consistency between frames. We also design a Bidirectional Spatial Mamba (BSM) block to bidirectionally process latent poses, to enhance accurate motion generation within a temporal frame. Our proposed method achieves up to 50% FID improvement and up to 4 times faster on the HumanML3D and KIT-ML datasets compared to the previous best diffusion-based method, which demonstrates strong capabilities of high-quality long sequence motion modeling and real-time human motion generation.
@@ 1 Introduction
Human motion generation involves creating high-quality 3D human motion based on given conditions, such as a text prompt, which generates Euler angles for each joint in a human skeleton. To emulate human motion effectively, virtual characters must respond to the conditional context, exhibit natural movement, and perform motion accurately. Recent motion generation models are categorized into four main approaches: autoencoder-based [1, 49, 37, 17, 59, 10], utilizing transformers for latent space compression and motion synthesis; GAN-based [31, 20, 4], using discriminators to enhance the realism of generated motions; autoregressive models [25], treating motion sequences as languages with specialized codebooks; and diffusion-based [55, 50, 6], employing denoising steps for motion generation. Challenges vary across methods, with autoencoder models struggling to generate accurate motions from detailed descriptions due to textual information compression, GAN-based models facing training difficulties, especially in conditional tasks, and diffusion-based models relying on complex transformer-based architectures results in inefficient of motion prediction.
@@ 2 Related Works
Human Motion Generation. Generating human motion is a significant application of computer vision, essential for various applications like 3D modelling and robot manipulation. Recently, the predominant method of achieving human motion generation, known as the Text-to-Motion task, involves learning a common latent space for both language and motion.
DVGAN [31] creates the GAN [11] discriminator by densely validating at each time-scale and perturbing the discriminator input for translation invariance, enabling motion generation and completion. ERD-QV [20] enhances latent representations through two additive modifiers: a time-to-arrival embedding applied universally and an additive scheduled target noise vector used during extended transitions. It further improves transition quality by incorporating a GAN framework with two discriminators operating at different timescales. HP-GAN [4], trained with a modified version of the improved WGAN-GP [16], utilizes a custom loss function designed for human motion prediction. It learns a probability density function of future human poses conditioned on previous poses.
Autoencoders [44, 28] are notable generative models known for their ability to represent data robustly by compressing high-dimensional data into a latent space, which is widely adopted in human motion generation [58, 57]. JL2P [1] uses RNN-based autoencoders [23] to learn a combined representation of language and pose. It restricts the direct mapping from text to motion to a one-to-one relationship. MotionCLIP [49] uses Transformer-based Autoencoders [52] to reconstruct motion while ensuring alignment with the corresponding text label in the CLIP [41] space. This alignment effectively integrates the semantic knowledge from CLIP into the human motion manifold. TEMOS [37] and T2M [17] combine a Transformer-based VAE [27] with a text encoder to generate distribution parameters that work within the VAE latent space. AttT2M [59] and TM2D [10] incorporate a body-part spatio-temporal encoder into VQ-VAE [51] for enhanced learning of a discrete latent space with increased expressiveness.
@@
State Space Models. Recently, state space sequence models (SSMs) [14, 15], drawing inspiration from classical state-space models [26], have emerged as a promising architecture for sequence modeling [24]. Mamba [12] introduces a selective SSM architecture, integrating time-varying parameters into the SSM framework, and proposes a hardware-aware algorithm to facilitate highly efficient training and inference processes. Some research works leverage SSM in computer vision to process 2D data. The 2D SSM [3] introduces an SSM block at the beginning of each transformer block [52, 8]. This approach aims to achieve efficient parameterization, accelerated computation, and a suitable normalization scheme. SGConvNeXt [30] presents a structured global convolution method inspired by ConvNeXt [33], incorporating multi-scale sub-kernels to achieve both parameterization efficiency and effective long sequence modeling. ConvSSM [47] integrates the tensor modeling principles of ConvLSTM [46] with SSMs, elucidating the utilization of parallel scans in convolutional recurrences. This approach enables subquadratic parallelization and rapid autoregressive generation. Vim [60] introduces a bidirectional SSM block [53] for efficient and versatile visual representation learning, achieving performance comparable to established ViT [8] methods. VMamba [32] introduces a Cross-Scan Module (CSM) designed to traverse the spatial domain and transform any non-causal visual image into ordered patch sequences. This approach achieves linear complexity while preserving global receptive fields. There have also been attempts to utilize SSMs to handle higher-dimensional data. Mamba-ND [29] explores various combinations of SSM and different scan directions within the SSM block to adapt Mamba [12] to higher-dimensional tasks. Recent efforts have sought to replace the traditional transformer-based U-Net within the diffusion denoiser with the SSM block, with the aim of enhancing image generation efficiency. DiffuSSM [54] adeptly manages higher resolutions without relying on global compression, thus maintaining detailed image representation throughout the diffusion process.
@@ 3.1 Preliminaries
Selective Structured State Space Sequence Model. SSMs particularly through the contributions of structured state space sequence models (S4) and Mamba, have demonstrated exceptional proficiency in handling long sequences. These models operationalize the mapping of a 1-D function or sequence, $x(t)\in\mathbb{R}\mapsto y(t)\in\mathbb{R}$, through a hidden state $h(t)\in\mathbb{R}^{N}$, employing $\mathbf{A}\in\mathbb{R}^{N\times N}$ as the evolution parameters, $\mathbf{B}\in\mathbb{R}^{N\times 1}$ and $\mathbf{C}\in\mathbb{R}^{1\times N}$ as the projection parameters, respectively.
@@ 3.2 Motion Mamba
$K=\{S_{2N-1},S_{2(N-1)-1},\ldots,S_{1}\}.$ (4)
This sequence specifies the number of scans allocated to each layer, in descending order of complexity. For instance, the uppermost encoder layer, $E_{1}$, and the lowermost decoder layer, $D_{N}$, are allocated $S_{2N-1}$ scans, indicating the highest scanning complexity. Conversely, the lowest encoder layer, $E_{N}$, and the uppermost decoder layer, $D_{1}$, are assigned $S_{1}$ scans, reflecting the lowest level of scanning complexity.
$E_{i}(S)=\begin{cases}S_{2N-1}&\text{for }i=1\\ S_{2(N-i)-1}&\text{for }i=2,\ldots,N-1\\ S_{1}&\text{for }i=N\end{cases}$ (5)
$D_{j}(S)=\begin{cases}S_{2N-1}&\text{for }j=N\\ S_{2(N-j)-1}&\text{for }j=N-1,\ldots,2\\ S_{1}&\text{for }j=1\end{cases}$ (6)
@@
Hierarchical Temporal Mamba (HTM) block processes compressed latent representations, denoted as $z$, with the dimensions $(T,B,C)$, of which procedure shown in Algorithm 1. Here, $T$ signifies the temporal dimension, as specified in the Variational AutoEncoder (VAE) framework. Initially, the input $z$ is subjected to a linear projection layer, producing transformed representations $x$ and $z$ with dimension $E$. Our analysis revealed an increased density of motion within the lower-level feature spaces. Consequently, we developed a hierarchical scanning methodology that is executed at various depths of the network. This approach not only accommodates the diverse motion densities, but also significantly reduces computational overhead. This step utilizes a hierarchically structured set of scans, $K=\{S_{2N_{n}-1},S_{2N_{n-1}-1},\ldots,S_{1}\}$, in conjunction with a corresponding series of memory matrices $\{A_{1},\ldots,A_{k}\}$. Each sub-SSM scan first applies a 1-D convolution to $x$, resulting in $x^{\prime}_{o}$. $x^{\prime}_{o}$ is then linearly projected to derive $B_{o}$, $C_{o}$, and $\Delta_{o}$. These projections $B_{o}$, $C_{o}$ use $\Delta_{o}$ to effect transformations in $\overline{A}_{o}$ and $\overline{B}_{o}$, respectively. After executing a sequence of SSM scans $\{SSM_{A_{1},x},SSM_{A_{2},x},\ldots,SSM_{A_{k},x}\}$, a set of outputs $\{O_{1},\ldots,O_{k}\}$ is compiled. This collection is subsequently aggregated via a linear projection to obtain the final output of the HTM block.
@@ 4.3 Comparative Studies
We evaluate our method against the state-of-the-art methods on the HumanML3D [17] and KIT-ML [39] datasets. We train our Motion Mamba with HTM arrangement strategy MM ($\{S_{2N_{n}-1},\ldots,S_{1}\}$), BSM bidirectional block strategy on the latent dimension = 2 with 11 layers. We evaluate our model and previous works with suggested metrics in HumanML3D [17] and calculate 95% confidence interval by repeat evaluation 20 times. The results for the HumanML3D dataset are presented in Table 1. Our model outperforms other methods significantly across various evaluation metrics, including FID, R precision, multi-modal distance, and diversity. For instance, our Motion Mamba outperforms previous best diffusion based motion generation MLD by 40.5% in terms of FID, and up to 10% improvement on R Precision, we aslo obatined best MModality by 3.060. The results for the KIT-ML dataset are presented in Table 2. We have also outperformed other well-established methods in FID and multi-modal distance.
@@ 4.4 Ablation Studies
Long Sequence Motion Generation. The HumanML3D [17] dataset exhibits a long-tailed and right-skewed distribution with a significant proportion of long-sequence human motions, as shown in Figure 3. We suggest previous studies overlooked the challenges in the long-sequence generation problem. Thus, we introduce a new dataset variant, HumanML3D-LS, comprising motion sequences longer than 190 frames extracted from the original test set. This addition allows us to showcase our capability in generating long-sequence motions. Subsequently, we evaluate the performance of our method on HumanML3D-LS and compare it with other diffusion-based motion generation approaches. The comparative results are presented in Table 3. Motion Mamba by leverage the benefits on long-range dependency modeling make it well suitable for long sequence motion generation.
Hierarchical Design with HTM. In our ablation studies, we observed a slight improvement upon reversing the scan order from a lower to a higher level, specifically transitioning from MM $\{S_{1},\ldots,S_{N}\}$ to MM $\{S_{N},\ldots,S_{1}\}$. This enhancement suggests a correlation with the increase in temporal motion density within the lower-level feature spaces. Furthermore, to achieve the optimal result, we introduce the hierarchical design to arrange the scanning frequency, resulting in the sequence MM $\{S_{2N_{n}-1},\ldots,S_{1}\}$. This expansion in the number of scans led to a performance increase. We attribute this enhancement to the observation that individual selective scan operations significantly reduce the parameter count, especially when compared to the parameter-intensive constructs of self-attention and feedforward network blocks prevalent in transformer architectures.
@@
MM ($\{S_{1},\ldots,S_{N}\}$) $0.673^{\pm.003}$ $1.278^{\pm.012}$ $3.802^{\pm.041}$ $8.678^{\pm.096}$ $3.127^{\pm.024}$
MM ($\{S_{N},\ldots,S_{1}\}$) $0.738^{\pm.002}$ $0.962^{\pm.011}$ $3.433^{\pm.003}$ $9.180^{\pm.071}$ $2.723^{\pm.033}$
MM ($\{S_{1},\ldots,S_{2N_{n-1}}\}$) $0.698^{\pm.002}$ $0.856^{\pm.008}$ $3.624^{\pm.037}$ $9.229^{\pm.067}$ $2.826^{\pm.017}$
MM ($\{S_{2N_{n}-1},\ldots,S_{1}\}$) $0.792^{\pm.002}$ $0.281^{\pm.009}$ $3.060^{\pm.058}$ $9.871^{\pm.084}$ $2.294^{\pm.058}$
@@ 5 Discussion and Conclusion
Acknowledgements
This work is partially supported by the Fundamental Research Funds for the Central Universities, Peking University.